OpenStudy (anonymous):

Please Help ASAP!? What is the solution of the equation? 3^5√(x+2)^3+3=27

OpenStudy (anonymous):

subtract 3, divide by 3^5, raise both sides to the 1/3 power, raise both sides to the 2 power, subtract 2.

OpenStudy (anonymous):

This is assuming that your sqrt looks like this (sqrt(x+2))^3

OpenStudy (anonymous):

\[3^5\sqrt{(x+2)^{3}}+3=27\] Divide by 3 to both sides. \[3^4\sqrt{(x+2)^{3}}+1=9\] \[3^4\sqrt{(x+2)^{3}}=9-1\] \[\sqrt{(x+2)^{3}}=\frac{ 8 }{ 3^{4} }\] \[\sqrt{(x+2)^{3}}=\frac{ 8 }{ 81 }\] \[(x+2)^{3}=\frac{ 64 }{ 6561 }\] \[x+2=\frac{ 4 }{ \sqrt[3]{6561} }\] \[x=\frac{ 4 }{ \sqrt[3]{6561} }-2\] \[x=\frac{ 4-2\sqrt[3]{6561} }{ \sqrt[3]{6561} }\]

OpenStudy (anonymous):

so x=4-2^3√6561/^3√6561 is my answer?

OpenStudy (janewelcomes):

@sophiasmommy94 was that the right answer?

OpenStudy (anonymous):

I hope so